DIFFERENTIAL GEOMETRY/PDE SEMINAR

Thursday, March 24, 2005 Padelford C-401 2:30–3:30 pm (Note: special day, time, and place!)

Local Isometric Embedding and Tricomi Equations

Qing Han (U of Notre Dame)

In this talk, I will give a simple proof of the following result: Any 2-dimensional Riemannian metric in a neighborhood of the origin in the plane admits a local isometric embedding in 3-space if Gauss curvature changes its sign cleanly at the origin. The proof is based on a simple construction of solutions to Tricomi equations.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.